Project

General

Profile

09152015

Nightly Observing Plan for MJD 57281 (starting evening of Sep 15, 2015)

DES has the full night.

Observers

  • Ofer Lahav (observer 1)
  • Devon Hollowood (observer 2)
  • Alex Kim (run manager)
  • Manuel Garcia-Fernandez (acclimating future observer)

CTIO Staff

  • Hernan Tirado - Obs. Support
  • Alberto Alvarez - operator
  • Staff calendar

Almanac

mjd 57281, date 20150915, Site CTIO, UT-offset 3 hr
Current local time 12:38 CLST     UT 15:38     SidTim 10:33
Sunset     19:42 CLST  22:42 UT   SidTim  264 deg  
Eve. -10   20:17 CLST  23:17 UT   SidTim  273 deg  Zeropoint; standards
Eve. -14   20:35 CLST  23:35 UT   SidTim  278 deg  Begin obstac
Eve. -18   20:54 CLST  23:54 UT   SidTim  282 deg  Last chance standards
Moonset    21:46 CLST  00:46 UT   SidTim  296 deg  
Midpoint   01:38 CLST  04:38 UT   SidTim  353 deg  
Morn -18   06:21 CLST  09:21 UT   SidTim   64 deg  Begin astronomical twilight
Morn Std   06:39 CLST  09:39 UT   SidTim   68 deg  End obstac; begin standards
Morn -10   06:59 CLST  09:59 UT   SidTim   74 deg  
Sunrise    07:33 CLST  10:33 UT   SidTim   82 deg  
Moon position (deg): RA 203.1, Dec -7.0 Phase: 0.151

Weather Conditions

Observing Plan and Goals for the Night

  • DES 20:17 CLT to 06:59 CLT
  • Beginning of night standards [Yzirg ordered scripts]
UT Date = '2015/09/15 23:17'     Local time = 20:17
SDSSJ2140-0000        ra 21:40:00  dec +00:00:00  secz 1.9
E6-A                  ra 14:45:33  dec -45:15:34  secz 1.4
E8-A                  ra 20:07:22  dec -44:37:01  secz 1.1
  • Obstac
  • morning standards [grizY ordered scripts]
UT Date = '2015/09/16 9:39'     Local time = 06:39
SDSSJ0100-0000        ra 01:00:00  dec +00:00:00  secz 2.0
SDSSJ0200-0000        ra 02:00:00  dec +00:00:00  secz 1.5
SDSSJ0320-0000        ra 03:20:00  dec +00:00:00  secz 1.2

Useful links

Nightsum | CTIO Night Report | DECam Logbook
Tololo Environmental Web Page | Detailed hourly forecast | Satelite maps

Notes from the 4PM Meeting

  •  Camera and Telescope - No problems.
  •  Last Night - Dome was not opened. Did DECal.
  • DESDM & DQ - Will look at this after phone con.
    *  Forecast/Plan for Tonight - ObsTac run. SN X field later in the night.
  • New ObsTac installed: Tweak 1) 20 degree moon hard cut fixed. Tweak 2) redder bands during dark time reduced.
  • Nightly calibration. Eric will upload new SISPI. We will ask Hernan to relaunch SISPI.

Other Notes

  • When were the SN fields last observered?
    hex_name    seqtype    last_complete
    SN-X3    long_i    2015-09-08 08:12:34.924511+00:00
    SN-X3    long_z    2015-09-08 08:46:41.064389+00:00
    SN-C3    long_r    2015-09-11 07:46:29.777327+00:00
    SN-S1    short    2015-09-11 09:09:54.635815+00:00
    SN-E1    short    2015-09-12 02:00:46.752304+00:00
    SN-S2    short    2015-09-12 07:22:38.092039+00:00
    SN-C3    long_z    2015-09-12 07:43:10.006001+00:00
    SN-C3    long_i    2015-09-12 08:50:47.646672+00:00
    SN-C2    short    2015-09-12 09:25:00.543599+00:00
    SN-E2    short    2015-09-13 04:26:58.318843+00:00
    SN-X1    short    2015-09-13 04:52:28.072487+00:00
    SN-X3    long_r    2015-09-13 05:12:03.453772+00:00
    SN-C1    short    2015-09-13 05:36:05.207316+00:00
    SN-C3    long_g    2015-09-13 05:55:43.368258+00:00
    SN-X3    long_g    2015-09-13 06:10:54.455087+00:00
    SN-X2    short    2015-09-13 07:09:42.924741+00:00